OpenStudy (katie_rogers101):

If there is a goat tied to a rectangular barn on a 50 foot lead and the barn is 20 feet by 20 feet (floor), what is the maximum grazing area? If there are regions you can't find the area of, provide as good an estimate as you can. Assume the goat is tied to a corner outside the barn, cannot get in, and that the barn is not grazing area. (Remember, this will be based on parts of circles, no other shapes...the goat's rope will only get shorter when he tries to go around the barn...) @jhonyy9

OpenStudy (katie_rogers101):

7. When the rope goes around the barn, what is the new radius? How much of a circle can it make without hitting the barn or overlapping area you've already found? What is that area? The new radius is 30. It can go 20 feet without overlapping. The area is (1/4) x 900 x pi = 225pi. 8. When the rope goes around the barn the other way, what is the new radius? How much of a circle can it make without hitting the barn or overlapping area you've already found? What is that area? it is the same as number seven because it is just a different 4th of the circle. So the radius is still 30. It can still go 20 feet. and the area is (1/4) x 900 x pi= 225 ***these answers are correct. my teacher has already reviewed them but you will need to look at them to answer #9***

OpenStudy (katie_rogers101):

9. The areas you found in 7 and 8 overlap each other. How much do they overlap? What *approximate* shape do they make? What is that area?

OpenStudy (katie_rogers101):

10. What is the total grazing area the goat can reach?

OpenStudy (katie_rogers101):

http://www.mathisfunforum.com/showimage.php?pid=288970&filename=barn.gif&preview=true Heres a picture that will help.

OpenStudy (katie_rogers101):

@jhonyy9 @xapproachesinfinity

OpenStudy (katie_rogers101):

@Alknowsall

OpenStudy (katie_rogers101):

@satellite73

OpenStudy (anonymous):

man this is a very com;licated problem lol ;)

jimthompson5910 (jim_thompson5910):

This is a pretty tricky problem, but I would set up a coordinate system. There's no set way to do it but one way is to place the origin at the lower right corner where the goat is tied up. So the goat is tied up at (0,0) which is one corner of the barn Based on the image, the other points are (0,20) -- top right corner (-20,20) -- top left corner (-20,0) -- bottom left corner

jimthompson5910 (jim_thompson5910):

after you have a coordinate system set up, you can find the equations of each circle why do this? because you'll need to figure out where the circles intersect. Using algebra is one way to do that

OpenStudy (anonymous):

yeah but dont u have to use some sort of equation idk which one exactly tho :(

jimthompson5910 (jim_thompson5910):

the yellow circle (in the image) would be (x-0)^2 + (y-0)^2 = 50^2 x^2 + y^2 = 2500 since it's centered at (0,0) and has radius of 50

OpenStudy (katie_rogers101):

here let me post the other questions to see if the information in them helps you

OpenStudy (anonymous):

to find the area of this main circle in the link uv provided you should use this equation ---> A=pie*r2

OpenStudy (katie_rogers101):

6. How much of the 50 foot circle can the goat reach without getting interrupted by the barn? What is that area? 3/4x 50^2 xpi= 1875 pi

jimthompson5910 (jim_thompson5910):

the green and red circles have the same radius (30) just different centers the red circle is centered at (-20,0) so (x+20)^2 + y^2 = 900 is the equation of the red circle

jimthompson5910 (jim_thompson5910):

x^2 + (y-20)^2 = 900 is the equation of the green circle

OpenStudy (anonymous):

3/4x 50^2 xpi= 1875 pi for this equation i think x=7.39 but im not sure tho

OpenStudy (retireed):

Fairly accurate drawing

OpenStudy (katie_rogers101):

I am still very confused

jimthompson5910 (jim_thompson5910):

does the last picture @retirEEd posted help at all? those circles represent where the goat can go (the max distance). Of course, because the barn is in the way, the goat can't go everywhere along each circle. It's more like in this drawing http://www.mathisfunforum.com/showimage.php?pid=288970&filename=barn.gif&preview=true but the equations help you figure out where the circles cross so you can find the approximate area of the overlapping region

OpenStudy (katie_rogers101):

@inkyvoyd maybe you can help too?

jimthompson5910 (jim_thompson5910):

ok let's go through this one step at a time here is the barn with corners A,B,C,D A = (0, 0) B = (0, 20) C = (-20, 20) D = (-20, 0) see attached I'm making point A the point where the goat is tied up

jimthompson5910 (jim_thompson5910):

if the goat is tied up at point A, and the rope is 50 ft, then it would be able to roam anywhere within this black circle (see attached)

jimthompson5910 (jim_thompson5910):

however, the walls of the barn get in the way, so in reality, the goat can trace out this 3/4 circle in red

jimthompson5910 (jim_thompson5910):

when the goat reaches the positive y axis, the rope will start to wrap around point B. If the goat keeps going to the left of the positive y axis, then you get this quarter circle in green

jimthompson5910 (jim_thompson5910):

repeat these steps but now go clockwise (instead of counterclockwise) and you get this quarter circle in blue

OpenStudy (katie_rogers101):

@Anaise how so?

OpenStudy (katie_rogers101):

also is there like a equation to set up that I can you to see this in numbers?

OpenStudy (katie_rogers101):

I am just all around confused by this. I know it should be simple but I feel like I'm making it harder than it is

jimthompson5910 (jim_thompson5910):

the area you want to find out is the area of the two overlapping regions (formed by the blue and green quarter circles) the yellow region shown

jimthompson5910 (jim_thompson5910):

notice how the corner points of this yellow region are C, E, F, G The coordinates of the points are C = (-20, 20) E = (-30, 20) F = (-20, 30) G = (-28.71, 28.71) point G is found by solving the system of equations (x+20)^2 + y^2 = 900 x^2 + (y-20)^2 = 900

jimthompson5910 (jim_thompson5910):

now in #9, they're asking for the approximate area so it sounds like we can sorta take a shortcut and form the kite CEGF, then use the area of a kite formula to find the approximate area. This isn't perfect of course, but it's an approximation so it should work on some level

OpenStudy (anonymous):

You can ignore the circle only advice and draw the radii from DG and BG and have triangles to find the areas of to get an exact result.

OpenStudy (anonymous):

It's more work though... aprox. is close and prob. better.

OpenStudy (mathmale):

Katie: Next time, PLEASE sketch (or present) the situation described in the problem right up front. I'd bet much info about this would become apparent from a sketch. It seems as though you've left all the sketching to other people to do. What are the limitations on the areas which the goat can reach? That's a relatively easy question, given that the goat is tied up with a 50-foot rope.

OpenStudy (anonymous):

so did you guys got the problem solved?

OpenStudy (retireed):

I can see how to, but after the question closed gave up. It is essentially three areas added... Area1 = (3/4)(pi)(50^2) three quarters of the large radius Area2 = (1/4)(pi)(30^2) one quarter of the a medium radius circle Area3 = (1/4)(pi)(30^2) one quarter of the other medium radius circle together minus an OVERLAP half circle Area4 = (1/2)(pi)(10^2) one half of the a medium radius circle minus approximately half a square where the two medium radius circles OVERLAP If you look up the post to my fairly accurate drawing attachment you will the last overlap. I am not smart enough to calculate that area, but it looks triangular.